### v1

@@ 4.3 Gradient Descent with Adaptive Step Size
  for $i=1,...,\infty$ do
     1. Compute the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ based on Theorem 1 or 3 to achieve the accuracy stated in Theorem 4.3.
     2. Use Algorithm 1 or 3 to estimate the gradient ^$\hat{\nabla}C(\hat{K}_{i})$.
      3. Update the gradient as $\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i})$, with $\eta_{i}\leq h_{\mathrm{GD}}(C(\hat{K}_{i}))$.
@@
To achieve the desired accuracy of the estimates stated in Theorem 4.3, the corresponding values of $l_{i},r_{i},n_{i}$, as defined in Algorithm 4, should be selected based on Theorem 1. Theorem 4.3 ensures that the cost will converge to the optimal solution with a predefined accuracy gap $\epsilon$, meaning $C(K^{*})+\epsilon$ will be reached with a certain probability. However, due to the estimation error in the gradient, further improvements in cost stop when the realized cost $C(\hat{K}_{i})\leq C(K^{*})+\epsilon$. To achieve higher accuracy, it is necessary to minimize the gradient estimation error. Nevertheless, the algorithm can only converge to the optimal when the gradient estimation is exact, which is not possible with a finite number of rollouts. The interpretation of the $\sigma$ is given in Remark 5. Regarding the step size $\eta_{i}$ stated in Theorem 4.3, from the expression of $h_{\mathrm{GD}}$ given in (62), we observe that as the noise level $\Sigma_{w}$ increases, the step size $\eta_{i}$ decreases, reflecting the observation analyzed in Section 3.4.
In Algorithm 4, the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ are determined adaptively to the cost $C(\hat{K}_{i})$, i.e., they are computed online inside the for-loop. This design is aligned with Theorem 4.3, which shows that the cost function $C(\hat{K}_{i})$ decreases probabilistically with each iteration. As the cost $C(\hat{K}_{i})$ decreases, this leads to a reduction in the required rollouts $n_{i}$ and rollout length $l_{i}$, while the exploration radius $r_{i}$ increases. Alternatively, the algorithm can be modified to compute these parameters offline, where for all iterations $i$, $n_{i},l_{i},r_{i}$ are calculated once based on the initial cost $C(\hat{K}_{0})$.
4.4 Natural Policy Gradient with Adaptive Step Size
Analogous to the model-free gradient descent method, we now turn to analyzing the convergence of the natural policy gradient method. {Theorem}[Model-free Natural Policy Gradient with Adaptive Step Size] Suppose the initial $\hat{K}_{0}\in\mathcal{S}$, and consider natural policy gradient with adaptive step size:
(44) $\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$
where ^$\hat{\nabla}C(\hat{K}_{i})$ and $\hat{\Sigma}_{\hat{K}_{i}}$ are the gradient and covariance estimates from Algorithm 1 and $0<\eta_{i}\leq\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert C(\hat{K}_{i})}{\lambda_{1}(\Sigma_{w})}}$. Given any accuracy $\epsilon>0$, probability $\delta\in(0,1)$ and $\sigma\in(0,1)$, if the estimation error of the gradient ^$\hat{\nabla}C(\hat{K}_{i})$ and the covariance $\hat{\Sigma}_{\hat{K}_{i}}$ satisfy:
(45a) $\mathbb{P}$^$\left\{\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}\right\}\geq\sqrt{\delta},$
(45b) $\mathbb{P}$$\left\{\lVert\hat{\Sigma}_{\hat{K}_{i}}-{\Sigma}_{\hat{K}_{i}}\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\sqrt{b_{\nabla}(C(\hat{K}_{i}))}}\right\}\geq\sqrt{\delta},$
where $h_{\nabla}$ was introduced in Lemma 2.2 and $b_{\nabla}$ was defined in (56) in Appendix A.5, then for any $C(\hat{K}_{i})\geq C(K^{*})+\epsilon$, the following inequality holds:
$\mathbb{P}\left\{C(\hat{K}_{i+1})-C(K^{*})\leq\kappa_{i}(C(\hat{K}_{i})-C(K^{*}))\Big{|}C(\hat{K}_{i})\right\}\geq\delta,$
@@ 4.4 Natural Policy Gradient with Adaptive Step Size
$\mathbb{P}\left\{\min_{i\in[0,n_{\mathrm{NPG}}]}C(\hat{K}_{i})-C(K^{*})\leq\epsilon\right\}\geq\delta^{n_{\mathrm{NPG}}},$
if the number of iterations $n_{\mathrm{NPG}}$ satisfies:
$n_{\mathrm{NPG}}\geq\frac{\lVert\Sigma_{K^{*}}\rVert}{2(1-\sigma)\eta_{\mathrm{NPG}}\lambda_{1}(\Sigma_{w})}\log\frac{C(K_{0})-C(K^{*})}{\epsilon},$
where ${\eta_{\mathrm{NPG}}}:=\min\limits_{i\in[0,n_{\mathrm{NPG}}]}\eta_{i}$. From Theorem 1, we can select the values of $l^{\nabla}_{i},r^{\nabla}_{i},n^{\nabla}_{i}$ to satisfy the requirement for the gradient estimation error, as indicated in (45a). Similarly, we can select $l^{\Sigma}_{i},r^{\Sigma}_{i},n^{\Sigma}_{i}$ to satisfy the requirement for the covariance estimation error, as described in (45b). To ensure both requirements are met, we can then set: $n_{i}\geq\max\{n^{\nabla}_{i},n^{\Sigma}_{i}\},l_{i}\geq\max\{l^{\nabla}_{i},l^{\Sigma}_{i}\},r_{i}\leq\min\{r^{\nabla}_{i},r^{\Sigma}_{i}\}$. These choices guarantee that the estimates of both the gradient and the covariance matrix satisfy their respective accuracy requirements. This combined selection of parameters is implemented in Algorithm 5 to ensure convergence with the desired accuracy and probability.
  for $i=0,...,\infty$ do
     1. Compute the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ based on Theorem 1 to achieve the accuracy stated in Theorem 4.4.
     2. Use Algorithm 1 to estimate the gradient ^$\hat{\nabla}C(\hat{K}_{i})$ and covariance $\hat{\Sigma}_{\hat{K}_{i}}$.
  end for
{Remark}
[Qualitative effect of gradient and covariance errors on convergence] Model-free algorithms are extensions of model-based policy gradient methods. However, in the model-free setting, the convergence rate is typically slower, reduced by a factor of ($1-\sigma$) compared to the model-based case. This reduction is due to estimation errors in the gradient and covariance matrices, which are inherent in model-free methods as they rely on data-driven estimates rather than exact model information. When the estimated gradient closely matches the true gradient, the model-free policy gradient approaches the same convergence rate as the model-based policy gradient, corresponding to $\sigma\to 0$. However, if the estimation error becomes large, such that $\sigma\to 1$, further policy improvements become impossible, and the algorithm can no longer guarantee convergence to the optimal with the desired accuracy. This effect of gradient estimation error on the convergence rate and suboptimality gap aligns with our observations in Section 3.4, as illustrated by the black dotted line, blue dashed line, and green dash-dotted line in Figure 1.
@@ 4.5 Comparison of Model-free PG for Noise and Noise-free Case
In Algorithm 1, three key quantities, exploration radius $r$, rollout length $l$, and the required rollouts $n$, play a central role in determining the estimation errors on gradient and covariance estimation. The presence of stochastic noise acting on the system (1b) during data collection influences these quantities in the following ways:
@@
  • Rollout length: The rollout length $l$ is different for the noise-free case and the noise case. In both scenarios, the rollout length plays a crucial role in estimating the gradient and covariance, whose true values are defined over an infinite horizon. However, finite approximations must be used. In the noise-free case, as described in [10, Lemma 23], the rollout length is determined by the desired accuracy, model parameters, and initial covariance $\Sigma_{0}$. In the presence of noise, as discussed in Lemma C.1, the rollout length is influenced not only by $\Sigma_{0}$, but also by the noise covariance $\Sigma_{w}$. Therefore, the rollout length in the noisy case must account for both the system’s initial conditions and noise disturbances. These relationships are quantitatively explored in Lemma C.1 through (64) and (65).
  • Required rollouts: The most significant difference between the noisy and noise-free cases comes from using the matrix concentration inequality for the trajectories consisting of noisy data. In the noise-free case, the boundedness of the data is guaranteed by an upper bound on the initial state, i.e., $\lVert x_{0}^{(k)}\rVert\leq L_{0}$ (see [10, Lemmas 27, 29]). In contrast, in the noisy case, particularly with unbounded Gaussian noise, the state remains bounded only with a certain probability (Appendix C.4). This boundedness in probability is then used to apply matrix concentration inequalities, providing an upper bound on the covariance of the samples(Appendix C.5).
@@ 5 Simulation Results
In this subsection, we use numerical simulations  [^1] The Matlab codes used to generate these results (in Section 3.4 and Section 5) are accessible from the repository: https://github.com/col-tasas/2025-PGforLQRwithStochastic to demonstrate the benefits of incorporating the variance reduction technique and adaptive step sizes in model-free policy gradient algorithms and the overall effect of noise on them. The matrices $(A,B,Q,R)$ and $\hat{K}_{0}$ are the same as described in Section 3.4. The rollout length, exploration radius, and number of rollouts are set to $l=100,~{}r=0.05,~{}n=1000$, respectively.
5.1 Model-free GD with Variance Reduction and Adaptive Step Sizes
The update of the policy is according to the gradient descent method, whose update law was defined in (43). The step size $\eta$ is chosen based on the noise level: for $\Sigma_{w}=10^{-4}I$ is set to $100$, while for $\Sigma_{w}=10^{-2}I$, it is reduced to $1$, qualitatively in accordance with the bounds established in Theorem 4.3. For the variance reduction technique, the number of rollouts to estimate the baseline function, $n_{b}$, is set as $20$. We consider a time-varying step size to illustrate the behavior of the algorithm with an adaptive step size. For noise $\Sigma_{w}=10^{-2}I$, the adaptive stepsize is set as $\eta_{i}=1.08^{i}$, increasing over iterations. The step size continues to grow until the cost no longer decreases significantly (until $i=15$).
Figure 2: Model-free GD with/out Variance Reduction and with/out Adaptive Step Size
Figure 2 shows the relative suboptimality gap of the iterated policy $\hat{K}_{i}$ as a function of the iteration counter $i$. From it, we can draw a number of observations:
  • Suboptimal gap: As analyzed in Theorem 4.3, due to the estimation error in the gradient, the model-free gradient descent algorithm does not converge to the optimal solution but finds suboptimal policies.
  • Step size selection: When the noise level is higher, the step size must be decreased, consistent with the expression of $h_{\mathrm{GD}}$ (see (62)) introduced in Theorem 4.3. For adaptive step sizes, the algorithm exhibits a faster convergence rate compared to using fixed step sizes, highlighting the advantages of adaptive strategies.
  • Variance reduction impact: For the policy gradient method with variance reduction (green solid line and black dashed line), the gradient estimation error is significantly reduced compared to the method without variance reduction. This leads to a faster convergence rate, as evidenced by comparing the green solid line with the red dotted line and the black dashed line with the blue dot-dashed line. The performance improvement from variance reduction is more pronounced in high-noise scenarios, consistent with the analysis in Theorem 3.
@@ 5.2 Model-free NPG with Adaptive Step Sizes
The natural policy gradient algorithm updates the gain $\hat{K}_{i}$ according to the natural policy gradient method, whose update law was defined in (44). The step size is chosen as $\eta=0.1$ across all noise levels $\Sigma_{w}=10^{-4}I,10^{-2}I,10^{0}I$. Similar to Model-free GD, the time varying step size is set as $\eta_{i}=1.01^{i}\eta$, increasing over iterations. The step size continues to grow until the cost no longer decreases significantly (until $i=15$).
Figure 3: Model-free NPG with/out Adaptive Step Size
As shown in Figure 3, the model-free NPG algorithm achieves nearly the same accuracy even as the noise $\Sigma_{w}$ increases, provided that $l,r,n$ remain unchanged. With higher noise levels, the estimation errors of ^$\hat{\nabla}C(\hat{K}_{i})$ and $\hat{\Sigma}_{\hat{K}_{i}}$ increase. In our analysis in Theorems 1 and 31, these estimation errors are explicitly captured. Due to the update rule in the natural policy gradient method (see (44)), the estimated gradient is multiplied by the inverse of the estimated covariance. Interestingly, the errors in these two terms cancel out due to this inversion. If the error could be analyzed directly for the product ^$\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}$ rather than treating them separately (as done in Proof C.9), significantly stronger theoretical guarantees could be obtained. This observation highlights an important direction for future research. For adaptive step sizes, the algorithm demonstrates a faster convergence rate compared to using fixed step sizes, further supporting the potential benefit of adaptive strategies.

### v2

@@ 4.4 Natural Policy Gradient with Adaptive Step Size
[…]
  1. Compute the required rollouts $n_{i}$, exploration radius $r_{i}$ and rollout length $l_{i}$ based on Theorem 4 to achieve the accuracy stated in Theorem 8.
  2. Use Algorithm 1 to estimate the gradient $\hat{\nabla}C(\hat{K}_{i})$ and covariance $\hat{\Sigma}_{\hat{K}_{i}}$.
@@
Theorem 8 (Model-free Natural Policy Gradient with Adaptive Step Size)
Suppose the initial $\hat{K}_{0}\in\mathcal{S}$, and consider natural policy gradient with adaptive step size:
$\hat{K}_{i+1}=\hat{K}_{i}-\eta_{i}\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1},\quad\forall i\in\mathbb{Z}_{+},$ (44)
where $\hat{\nabla}C(\hat{K}_{i})$ and $\hat{\Sigma}_{\hat{K}_{i}}$ are the gradient and covariance estimates from Algorithm 1 and $0<\eta_{i}\leq\frac{1}{2\lVert R\rVert+\frac{2\lVert B\rVert C(\hat{K}_{i})}{\lambda_{1}(\Sigma_{w})}}$.
Given any accuracy $\epsilon>0$ and $\sigma\in(0,1)$, defining ${\eta_{\mathrm{NPG}}}:=\inf\limits_{i}\eta_{i}$ and the number of iterations $n_{\mathrm{NPG}}$:
$n_{\mathrm{NPG}}\geq\frac{\lVert\Sigma_{K^{*}}\rVert}{2(1-\sigma)\eta_{\mathrm{NPG}}\lambda_{1}(\Sigma_{w})}\log\frac{C(K_{0})-C(K^{*})}{\epsilon}.$
Given any probability $\delta\in(0,1)$ satisfying $\delta n_{\mathrm{NPG}}\in(0,1)$, assume that the estimation error of the gradient $\hat{\nabla}C(\hat{K}_{i})$ and the covariance $\hat{\Sigma}_{\hat{K}_{i}}$ satisfy:
$\mathbb{P}$ $\left\{\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}\right\}\geq\sqrt{1-\delta},$ (45a)
$\mathbb{P}$ $\left\{\lVert\hat{\Sigma}_{\hat{K}_{i}}-{\Sigma}_{\hat{K}_{i}}\rVert\leq\frac{\sigma\epsilon\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\sqrt{b_{\nabla}(C(\hat{K}_{i}))}}\right\}\geq\sqrt{1-\delta},$ (45b)
where $h_{\nabla}$ was introduced in Lemma 5 and $b_{\nabla}$ was defined in (56) in Appendix A.5. Then for any $C(\hat{K}_{i})\geq C(K^{*})+\epsilon$, the following inequality holds:
$\mathbb{P}\left\{C(\hat{K}_{i+1})-C(K^{*})\leq\kappa_{i}(C(\hat{K}_{i})-C(K^{*}))\Big{|}C(\hat{K}_{i})\right\}\geq 1-\delta,$
@@
$\mathbb{P}\left\{\min_{i\in[0,n_{\mathrm{NPG}}]}C(\hat{K}_{i})-C(K^{*})\leq\epsilon\right\}\geq 1-\delta{n_{\mathrm{NPG}}},$
From Theorem 4, we can select the values of $l^{\nabla}_{i},r^{\nabla}_{i},n^{\nabla}_{i}$ to satisfy the requirement for the gradient estimation error, as indicated in (45a). Similarly, we can select $l^{\Sigma}_{i},r^{\Sigma}_{i},n^{\Sigma}_{i}$ to satisfy the requirement for the covariance estimation error, as described in (45b). To ensure both requirements are met, we can then set: $n_{i}\geq\max\{n^{\nabla}_{i},n^{\Sigma}_{i}\},l_{i}\geq\max\{l^{\nabla}_{i},l^{\Sigma}_{i}\},r_{i}\leq\min\{r^{\nabla}_{i},r^{\Sigma}_{i}\}$. These choices guarantee that the estimates of both the gradient and the covariance matrix satisfy their respective accuracy requirements. This combined selection of parameters is implemented in Algorithm 5 to ensure convergence with the desired accuracy and probability.
Remark 3 (Qualitative effect of gradient and covariance errors on convergence)
Model-free algorithms are extensions of model-based policy gradient methods. However, in the model-free setting, the convergence rate is typically slower, reduced by a factor of ($1-\sigma$) compared to the model-based case. This reduction is due to estimation errors in the gradient and covariance matrices, which are inherent in model-free methods as they rely on data-driven estimates rather than exact model information. When the estimated gradient closely matches the true gradient, the model-free policy gradient approaches the same convergence rate as the model-based policy gradient, corresponding to $\sigma\to 0$. However, if the estimation error becomes large, such that $\sigma\to 1$, further policy improvements become impossible, and the algorithm can no longer guarantee convergence to the optimal with the desired accuracy. This effect of gradient estimation error on the convergence rate and suboptimality gap aligns with our observations in Section 3.4, as illustrated by the black dotted line, blue dashed line, and green dash-dotted line in Figure 1.
@@ 4.5 Comparison of Model-free PG for Noise and Noise-free Case
In Algorithm 1, three key quantities, i.e. exploration radius $r$, rollout length $l$, and the required rollouts $n$, play a central role in determining the estimation errors on gradient and covariance estimation. The presence of stochastic noise acting on the system (1b) during data collection influences these quantities in the following ways:
@@
  • Rollout length: The rollout length $l$ is different for the noise-free case and the noise case. In both scenarios, the rollout length plays a crucial role in estimating the gradient and covariance, whose true values are defined over an infinite horizon. However, finite approximations must be used. In the noise-free case, as described in [9, Lemma 23], the rollout length is determined by the desired accuracy, model parameters, and initial covariance $\Sigma_{0}$. In the presence of noise, as discussed in Lemma 6, the rollout length is influenced not only by $\Sigma_{0}$, but also by the noise covariance $\Sigma_{w}$. Therefore, the rollout length in the noisy case must account for both the system’s initial conditions and noise disturbances. These relationships are quantitatively explored in Lemma 6.
  • Required rollouts: The most significant difference between the noisy and noise-free cases comes from using the matrix concentration inequality for the trajectories consisting of noisy data. In the noise-free case, the boundedness of the data is guaranteed by an upper bound on the initial state, i.e., $\lVert x_{0}^{(k)}\rVert\leq L_{0}$ (see [9, Lemmas 27, 29]). In contrast, in the noisy case, particularly with unbounded Gaussian noise, the state remains bounded only with a certain probability (Appendix C.4). This boundedness in probability is then used to apply matrix concentration inequalities, providing an upper bound on the covariance of the samples(Appendix C.5 and [51, Appendix C.6]).
@@ 5 Simulation Results
In this subsection, we use numerical simulations  [^1] The Matlab codes used to generate these results (in Section 3.4 and Section 5) are accessible from the repository: https://github.com/col-tasas/2025-PGforLQRwithStochastic to demonstrate the benefits of incorporating the variance reduction technique and adaptive step sizes in model-free policy gradient algorithms and the overall effect of noise on them. The matrices $(A,B,Q,R)$ and $\hat{K}_{0}$ are the same as described in Section 3.4. The rollout length, exploration radius, and number of rollouts are set to l=100,r=0.04,n=1000l=100,~r=0.04,~n=1000, respectively. The simulation results are obtained from a Monte Carlo simulation over $5$ data samples.
5.1 Noise Level and Stepsize
The policy is updated using the policy gradient descent method in Algorithm 4. Figure 2 illustrates the relative suboptimality gap of the cost associated with the iterated policy $\hat{K}_{i}$ as a function of the iteration index $i$, for different noise levels and step sizes.
Figure 2: Effect of noise level on step size
From Figure 2, when the noise level is low, $\Sigma_{w}=10^{-4}I$, the step size is empirically chosen as $40$, as shown by the red dotted line. In this case, we observe a persistent suboptimality gap, which originates from the estimation error of the gradient inherent to the zeroth-order method from noisy data. When the noise level increases to $\Sigma_{w}=10^{-2}I$, a large step size ($\eta=6$) leads to divergence of the cost function (as shown in black dashed line). As stated in Theorem 7, when the noise level is higher, it is necessary to decrease the step size to ensure convergence. By reducing the step size to $\eta=0.3$, as illustrated by the blue solid line in the figure, the cost converges again with a suboptimality gap.
5.2 Effect of variance reduction on PGD
The step size $\eta$ is chosen based on the noise level: for $\Sigma_{w}=10^{-4}I$ it is set to $40$, while for $\Sigma_{w}=10^{-2}I$, it is reduced to $0.3$, qualitatively in accordance with the bounds established in Theorem 7. For the variance reduction technique, the number of rollouts to estimate the baseline function, $n_{b}$, is set to $200$.
Figure 3: PGD with/out variance reduction
Figure 3 illustrates the relative suboptimality gap of the cost associated with the iterated policy $\hat{K}_{i}$ as a function of the iteration index $i$. We compare two groups of curves corresponding to different noise levels: the blue dashed and black solid lines for the noise level $\Sigma_{w}=10^{-4}I$, and the green dash-dotted and magenta dotted lines for $\Sigma_{w}=10^{-2}I$. The blue dashed and green dash-dotted lines represent results without variance reduction. It can be observed that the performance improvement achieved through variance reduction is more pronounced in high-noise scenarios, which is consistent with the analysis in Theorem 6.
@@ 5.3 Model-free NPG with Adaptive Step Sizes
The policy is updated using the Natural policy gradient method in Algorithm 5. For the case of fixed step size it is chosen $\eta=\frac{a}{b+c\mathrm{Tr(P_{\hat{K}_{0}})}}$ which follows the structure of the step size derived in Theorem 8 and is applied uniformly across all noise levels $\Sigma_{w}={10^{-4}I,10^{-2}I,10^{0}I}$. The parameters $a$, $b$, and $c$ are selected empirically as $0.09$, 11, and $2$, respectively. For the case of adaptive step sizes, we take $\eta_{i}=\frac{a}{b+c\mathrm{Tr(P_{\hat{K}_{i}})}}$, where $\mathrm{Tr}(P_{K_{i}})$ reflects the current cost $C(\hat{K}_{i})$ associated with the policy at iteration $i$.
Figure 4: NPG with/out adaptive step size
Comparing the black solid line and the green dashed line in Figure 4, we observe that using adaptive step sizes yields a faster convergence rate than fixed step sizes, further highlighting the benefits of adaptive strategies. As shown in Figure 4 (red dotted, green dashed, and blue dot–dashed lines), the model-free NPG algorithm achieves nearly the same accuracy even as the noise $\Sigma_{w}$ increases, provided that $l,r,$ and $n$ remain unchanged. At higher noise levels, the estimation errors of both $\hat{\nabla}C(\hat{K}_{i})$ and $\hat{\Sigma}_{\hat{K}_{i}}$ grow. These errors are explicitly accounted for in our analysis (Theorems 4 and 5). Due to the natural policy gradient update rule (see (44)), the estimated gradient is multiplied by the inverse of the estimated covariance, which causes the errors in these two quantities to partially offset each other. If the analysis were performed directly on the product $\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}$ rather than treating them separately (as in Proof of Theorem 2), substantially stronger theoretical guarantees could be established. This observation points to an important direction for future research.